Sweatshirt Inc’s ROE is 20%. It’s dividend payout ratio is 70%. The last dividend, just
paid, was RM2.00. If its dividends are expected to grow at the company's internal
growth rate indefinitely, what is the current value of the company's common stock if
its required return is 18%?
1
Expert's answer
2015-03-23T11:03:47-0400
ROE = 20%, dividend payout ratio is 70%, required return is 18%. The last dividend was RM2.00. If its dividends are expected to grow at the company's internal growth rate indefinitely, the current value of the company's common stock is: P = div0/(r - g) = 2/(0.18 - 0.2*0.7) = RM50. The formula for the present value of a stock with constant growth is the estimated dividends to be paid divided by the difference between the required rate of return and the growth rate.